A certain fish requires 3 hours to swim 15 kilometers downstream. The return trip upstream take 5 hours. Find the speed of the fish in still water. How fast is the current of the stream?
Answer by Aswathy(23)

Let x and y be the speed of fish in still water and water current respectively.
Then the speed of fish
(1)when it swims downstream,
x+y=15/3=5 (v=s/t)
(2)when it swims upstream,
x-y=15/5=3
so by solving the above pair of linear equation we get x=4 & y=1.
So, the speed of the fish in still water is 4km/h & the speed of the current is 1km/h.